### v1

@@ 2.2.2 Second-order policy optimization
where $\boldsymbol{H}_{k}$ denotes the general Hessian matrix of the objective function with respect to its decision variables at the $k$-th iterate. It can be shown that this update rule delivers a quadratic convergence rate (Nocedal and Wright, 2006). In case of (9), $\boldsymbol{H}_{k}$ refers to the Hessian of the expected cumulative reward with respect to the policy parameters $\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})\in\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$ evaluated at the $k$-th iterate $\boldsymbol{H}_{k}=\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{k}}$. However, the computation of the exact deterministic policy Hessian is computationally expensive and therefore prohibitive for most applications.
Quasi-Newton methods are a computationally tractable alternative to exact Newton methods not only for RL but for optimization in general. The main idea of quasi-Newton methods is to approximate the exact Hessian with a computationally tractable alternative $\boldsymbol{B}_{k}\approx\boldsymbol{H}_{k}$ at the expense of a reduced superlinear convergence rate (Nocedal and Wright, 2006). In alignment to Newton methods, the exact Hessian is interchanged with the tractable approximation resulting in the following update rule for quasi-Newton methods
@@
With a focus on policy optimization via RL, Kordabad et al. (2022) propose a computationally cheaper approximation of the deterministic policy Hessian. The authors show that the following expression approximates the deterministic policy Hessian $\boldsymbol{M}_{k}=\boldsymbol{B}_{k}\approx\boldsymbol{H}_{k}$ such that still a superlinear convergence rate can be achieved
@@
In the context of RL, the policy Hessian approximation (20) is computed using an expected value like the policy gradient. This expected value must be estimated empirically with a finite amount of samples similar to (13), resulting in $\tilde{\boldsymbol{B}}_{k}\approx\boldsymbol{B}_{k}$. Subsampled Newton methods form the subcategory of Newton methods that consider optimization with noisy gradient $\tilde{\boldsymbol{g}}_{k}$ and Hessian estimates $\tilde{\boldsymbol{B}}_{k}$. Detailed reviews on subsampled Newton methods are provided in Bollapragada et al. (2019) and Bottou et al. (2018). Analogous to (14) and (19), the update rule for subsampled Newton methods is derived as
@@
One central idea in subsampled Newton methods is that the quality of the obtained update in (23) is mainly determined by the accuracy of the gradient estimate $\tilde{\boldsymbol{g}}_{k}$. The accuracy of the Hessian estimate only has a minor influence on the proposed update rule as long as it is of sufficient quality. In many applications the Hessian estimate can be improved by taking large enough samples sizes. In practical RL approaches however, the Hessian estimation is more challenging because the action-value function $Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})$ is typically approximated. Insufficient action-value function approximation can affect the Hessian estimate severely leading to unstable updates. In the context of general large-scale optimization, approaches like AdaHessian (Yao et al., 2021) or Sophia (Liu et al., 2024) aim to alleviate instabilities due to insufficient Hessian estimation with the use of momentum on the Hessian estimate. Both approaches constrain themselves to vectorial Hessian approximations as they are particularly designed for large scale optimization and do not consider a good initialization of the exponentially averaged Hessian estimate.
@@ 2.3 Parameterized model predictive control as policy in reinforcement learning
RL agents act according to their policy. Most works in the area of RL focus on approximating the optimal policy using linear functions or neural networks (NNs). Despite having a good interpretability, linear functions are often limited in their expressiveness. NNs on the other side can represent almost any desired policy but lack interpretability due to their black-box mapping. In addition to that, if the NN cannot be pretrained, the NN parameters are usually initialized randomly, leading to an excessive amount of required training data. In this work, we focus on the approximation of the optimal policy via a parameterized MPC scheme.
MPC is a model-based optimal control approach that computes a sequence of optimal control actions $\boldsymbol{\mathrm{u}}^{\star}=({\boldsymbol{u}_{0}^{\star}},\ldots,{\boldsymbol{u}_{N-1}^{\star}})$ with $\boldsymbol{u}_{i}\in\mathcal{A}$ along a prediction horizon $N\in\mathbb{N}$ based on an internal system model such that a control objective is optimized and process constraint are satisfied. Gros and Zanon (2020) show that the following parameterized MPC optimization problem can be used as a policy in RL to capture the optimal policy
$\boldsymbol{\mathrm{u}}^{\star}(\boldsymbol{s})=\arg\min_{\boldsymbol{\mathrm{u}}}\leavevmode\nobreak\ \leavevmode\nobreak$ $V_{\mathrm{f},\boldsymbol{\theta}}(\boldsymbol{x}_{N})+\boldsymbol{w}_{\mathrm{f}}^{\top}\boldsymbol{\sigma}_{N}$
@@
$\mathrm{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\boldsymbol{x}_{i+1}=f^{\mathrm{MPC}}_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{u}_{i}),\leavevmode\nobreak\ \boldsymbol{x}_{0}=\boldsymbol{s}$ (24b)
@@
$h_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})=0.$ (24e)
In (24), the objective function (24a) is minimized. It consists of the sum of stage costs $\ell_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ starting from the initial condition at $i=0$ up to the second to last point of the prediction horizon at $i=N-1$, the terminal cost at the end of the prediction $V_{\mathrm{f},\boldsymbol{\theta}}:\mathcal{S}\rightarrow\mathbb{R}$, and penalties for soft constraints. The system dynamics are accounted for via the system equations $f^{\mathrm{MPC}}_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ starting from the initial state $\boldsymbol{x}_{0}$ as stated in (24b). The predicted trajectory does not need to be the same as in the MDP since $f^{\mathrm{MPC}}_{\boldsymbol{\theta}}$ is usually only an approximation of the true system dynamics (1). Maximum and minimum values for the control action are posed via (24c). Process equality and inequality constraints are ensured via $h_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n_{h}}$ and $g_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n_{g}}$ as shown in (24e) and (24d) respectively. To always ensure feasibility, the inequality constraints need to be relaxed to soft constraints. All functions with the index $\boldsymbol{\theta}$ are parameterizable and can be adapted by the RL algorithm. Note that to distinguish the states and actions within the optimal control problem (24) from those within the MDP, we introduce the states and actions within the MPC as $\boldsymbol{x}$ and $\boldsymbol{u}$. After (24) is solved, the MPC policy is derived by taking the first element of the sequence of optimal control actions $\boldsymbol{\mathrm{u}}^{\star}$, so $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})=\boldsymbol{u}_{0}^{\star}$.
@@ 2.3.1 Parametric NLP sensitivities
In order to update the policy parameters with a policy optimization method, the Jacobian of the policy with respect to its parameters needs to be computed. When using MPC as a policy, the derivative of the policy equals the derivative of the solution of the optimization problem, which is only computable explicitly in rare cases. The MPC optimization problem (24) can be categorized as a parametric constrained nonlinear program (NLP). The derivative of the optimal solution with respect to the parameters of the NLP is also often referred to as the (parametric) NLP sensitivities. We will show next, how the NLP sensitivities can be computed.
@@
$\boldsymbol{z}^{\star}(\boldsymbol{\zeta})=\arg\min_{\boldsymbol{z}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\Phi(\boldsymbol{z};\boldsymbol{\zeta})$ (25a)
$\mathrm{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $g(\boldsymbol{z};\boldsymbol{\zeta})\leq 0$ (25b)
@@
where $\boldsymbol{z}\in\mathbb{R}^{n_{\boldsymbol{z}}}$ denotes the decision variables, $\boldsymbol{\zeta}\in\mathbb{R}^{n_{\boldsymbol{\zeta}}}$ denotes the changeable parameters, $\Phi:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}$ denotes the scalar objective function, $g:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}^{n_{g}}$ denotes the inequality constraints, and $h:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}^{n_{h}}$ denotes the equality constraints. The optimal solution $\boldsymbol{z}^{\star}(\boldsymbol{\zeta})$ of (25) changes with the optimization parameters $\boldsymbol{\zeta}$. Except for rare cases, this functional relationship is typically unknown. We introduce the Lagrangian $\mathcal{L}:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{g}}\times\mathbb{R}^{n_{h}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}$ of (25) as
@@
where $\boldsymbol{\lambda}\in\mathbb{R}^{n_{g}}$ and $\boldsymbol{\nu}\in\mathbb{R}^{n_{h}}$ are the Lagrange multipliers, also known as dual variables, associated to the inequality and equality constraints respectively. To simplify notation, we introduce the stacked primal-dual vector $\boldsymbol{\chi}^{\top}=(\boldsymbol{z}^{\top},\boldsymbol{\lambda}^{\top},\boldsymbol{\nu}^{\top})$ with $\boldsymbol{\chi}\in\mathbb{R}^{n_{\boldsymbol{\chi}}}$ and $n_{\boldsymbol{\chi}}=n_{\boldsymbol{z}}+n_{g}+n_{h}$. Let $\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})$ denote an optimal primal-dual vector solving (25) and satisfying the Karush-Kuhn-Tucker (KKT) conditions (Nocedal and Wright, 2006), and let $g$ satisfy the linear-independency constraint qualifications (Nocedal and Wright, 2006), then the equalities of the KKT conditions form an implicit function $F:\mathbb{R}^{n_{\boldsymbol{\chi}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}^{n_{\boldsymbol{\chi}}}$ of the form
@@
where the operator $\odot$ denotes the Hadamard (elementwise) product. In order to obtain the first-order NLP sensitivities of the optimal primal-dual solution with respect to the NLP parameters $\nabla_{\boldsymbol{\zeta}}\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})\in\mathbb{R}^{n_{\boldsymbol{\chi}}\times n_{\boldsymbol{\zeta}}}$, (27) can be implicitly differentiated almost everywhere (Fiacco and Ishizuka, 1990) forming the following linear system
@@
with the coefficient matrix $\nabla_{\boldsymbol{\chi}^{\star}}F(\boldsymbol{\chi}^{\star};\boldsymbol{\zeta})\in\mathbb{R}^{n_{\boldsymbol{\chi}}\times n_{\boldsymbol{\chi}}}$ and the right-hand-side matrix $-\nabla_{\boldsymbol{\zeta}}F(\boldsymbol{\chi}^{\star};\boldsymbol{\zeta})\in\mathbb{R}^{n_{\boldsymbol{\chi}}\times n_{\boldsymbol{\zeta}}}$. The first-order NLP sensitivities can therefore be obtained by solving (28) for $\nabla_{\boldsymbol{\zeta}}\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})$. In the context of MPC as policy in RL, Brandner and Lucia (2024) show that the second-order NLP sensitivities can be obtained in a similar fashion and also boil down to the solution of a linear system. However the right-hand-side matrix requires third partial derivatives of the Lagrangian with respect to full primal-dual vector. This can render the computation intractable, for instance if the number of primal-dual variables $\boldsymbol{\chi}$ is large, or if highly nonlinear terms appear in the NLP such that the third partial derivatives become extremely convoluted.
To compute the deterministic policy gradient (12) and the respective terms of the deterministic policy Hessian (20), the NLP (25) must be cast into (24) by setting $\boldsymbol{z}^{\star}=\boldsymbol{\mathrm{u}}^{\star}$ and $\boldsymbol{\zeta}=\boldsymbol{\theta}$. The relevant sensitivities $\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})=\nabla_{\boldsymbol{\theta}}\boldsymbol{u}_{0,\boldsymbol{\theta}}^{\star}(\boldsymbol{s})$ must then be extracted from the full sensitivity matrix.

### v2

@@ 2.2.2 Second-order policy optimization
where $\boldsymbol{H}_{k}$ denotes the general Hessian matrix of the objective function with respect to its decision variables at the $k$-th iterate. It can be shown that this update rule delivers a quadratic convergence rate (nocedalNumericalOptimization2006). In case of (9), $\boldsymbol{H}_{k}$ refers to the Hessian of the expected cumulative reward with respect to the policy parameters $\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})\in\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$ evaluated at the $k$-th iterate $\boldsymbol{H}_{k}=\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{k}}$. However, the computation of the exact deterministic policy Hessian is computationally expensive and therefore prohibitive for most applications.
Quasi-Newton methods are a computationally tractable alternative to exact Newton methods not only for RL but for optimization in general. The main idea of quasi-Newton methods is to approximate the exact Hessian with a computationally tractable alternative $\boldsymbol{B}_{k}\approx\boldsymbol{H}_{k}$ at the expense of a reduced superlinear convergence rate (nocedalNumericalOptimization2006). In alignment to Newton methods, the exact Hessian is interchanged with the tractable approximation resulting in the following update rule for quasi-Newton methods
@@
With a focus on policy optimization via RL, kordabadQuasiNewtonIterationDeterministic2022 propose a computationally cheaper approximation of the deterministic policy Hessian. The authors show that the following expression approximates the deterministic policy Hessian $\boldsymbol{M}_{k}=\boldsymbol{B}_{k}\approx\boldsymbol{H}_{k}$ such that still a superlinear convergence rate can be achieved
@@
In the context of RL, the policy Hessian approximation (22) is computed using an expected value like the policy gradient. This expected value must be estimated empirically with a finite amount of samples similar to (13), resulting in $\tilde{\boldsymbol{B}}_{k}\approx\boldsymbol{B}_{k}$. Subsampled Newton methods form the subcategory of Newton methods that consider optimization with noisy gradient $\tilde{\boldsymbol{g}}_{k}$ and Hessian estimates $\tilde{\boldsymbol{B}}_{k}$. Detailed reviews on subsampled Newton methods are provided in bollapragadaExactInexactSubsampled2019 and bottouOptimizationMethodsLargeScale2018. Analogous to (14) and (21), the update rule for subsampled Newton methods is derived as
@@
One central idea in subsampled Newton methods is that the quality of the obtained update in (25) is mainly determined by the accuracy of the gradient estimate $\tilde{\boldsymbol{g}}_{k}$. The accuracy of the Hessian estimate only has a minor influence on the proposed update rule as long as it is of sufficient quality. In many applications the Hessian estimate can be improved by taking large enough sample sizes. In the context of general large-scale optimization, approaches like AdaHessian (yaoADAHESSIANAdaptiveSecond2021) or Sophia (liuSophiaScalableStochastic2024) aim to alleviate instabilities due to insufficient Hessian estimation with the use of momentum on the Hessian estimate. Both approaches constrain themselves to vectorial Hessian approximations as they are particularly designed for large scale optimization and do not consider a good initialization of the exponentially averaged Hessian estimate.
Still, even if the sample size is large enough, the Hessian can be singular. Even if the Hessian is regularized before the Newton update is computed, the update direction can be severely distorted, suggesting large steps into the directions with local linear behavior. Trust-region approaches constrain the length of the computed update to a certain trust-region radius $\delta_{k}>0$, which can introduce a higher level of stability. The goal is then to find the optimal step within this trust-region. In order to compute the trust-region constrained update step $\boldsymbol{p}_{k}$ the following optimization problem must be solved (nocedalNumericalOptimization2006), which results from a quadratic approximation of the objective function around the current iterate $\boldsymbol{\theta}_{k}$
$\boldsymbol{p}_{k}=\arg\min_{\boldsymbol{p}}\quad$ $\boldsymbol{p}^{\top}\tilde{\boldsymbol{g}}_{k}+\frac{1}{2}\boldsymbol{p}^{\top}\tilde{\boldsymbol{B}}_{k}\boldsymbol{p}$ (26a)
$\mathrm{s.t.}\quad$ $\left\|\boldsymbol{p}\right\|_{2}\leq\delta_{k}.$ (26b)
The trust-region radius $\delta_{k}$ does not need to be constant and can be adapted depending on how well the objective function and its second-order approximation align. However, these comparisons are often impractical for stochastic objective functions as they are in RL because the objective estimates are noisy, rendering the comparison unreliable. Therefore, we will suggest a systematic approach to update the trust-region radius without comparison to the objective function.
@@ 2.3 Parameterized model predictive control as policy in reinforcement learning
RL agents act according to their policy. Most works in the area of RL focus on approximating the optimal policy using NNs nowadays. Despite having a good interpretability, linear functions are often limited in their expressiveness. NNs on the other side can represent almost any desired policy but lack interpretability due to their black-box mapping. In addition to that, if the NN cannot be pretrained using a reasonable behavior policy, the NN parameters are usually initialized randomly requiring an end-to-end RL process, leading to an excessive amount of required training data. In this work, we focus on the approximation of the optimal policy via a parameterized MPC scheme.
MPC is a model-based optimal control approach that computes a sequence of optimal control actions $\boldsymbol{\mathrm{u}}^{\star}=({\boldsymbol{u}_{0}^{\star}},\ldots,{\boldsymbol{u}_{N-1}^{\star}})$ with $\boldsymbol{u}_{i}\in\mathcal{A}$ along a prediction horizon $N\in\mathbb{N}$ based on an internal system model such that a control objective is optimized and process constraint are satisfied. grosDataDrivenEconomicNMPC2020 show that the following parameterized MPC optimization problem can be used as a policy in RL to capture the optimal policy
$\boldsymbol{\mathrm{u}}^{\star}(\boldsymbol{s})=\arg\min_{\boldsymbol{\mathrm{u}}}\penalty 10000\ \penalty 10000\$$V_{\mathrm{f},\boldsymbol{\theta}}(\boldsymbol{x}_{N})+\boldsymbol{w}_{\mathrm{f}}^{\top}\boldsymbol{\sigma}_{N}$
@@
$\mathrm{s.t.}\penalty 10000\$$\boldsymbol{x}_{i+1}=f^{\mathrm{MPC}}_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{u}_{i}),\penalty 10000\ \boldsymbol{x}_{0}=\boldsymbol{s}$ (27b)
@@
$h_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})=0$ (27e)
$g_{\mathrm{f},\boldsymbol{\theta}}(\boldsymbol{x}_{N})\leq\boldsymbol{\sigma}_{N},\quad\boldsymbol{\sigma}_{N}\geq 0.$ (27f)
In (27), the objective function (27a) is minimized. It consists of the sum of stage costs $\ell_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ starting from the initial condition $\boldsymbol{s}$ at $i=0$ up to the second to last point of the prediction horizon at i=N1i=N-1, the terminal cost at the end of the prediction $V_{\mathrm{f},\boldsymbol{\theta}}:\mathcal{S}\rightarrow\mathbb{R}$, and penalties for soft constraints. The system dynamics are accounted for via the system equations $f^{\mathrm{MPC}}_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ starting from the initial state $\boldsymbol{x}_{0}$ as stated in (27b). The predicted trajectory does not need to be the same as in the MDP since $f^{\mathrm{MPC}}_{\boldsymbol{\theta}}$ is usually only an approximation of the true system dynamics (1). Maximum and minimum values for the control action are posed via (27c). Process equality and inequality constraints are ensured via $h_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n_{h}}$ and $g_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n_{g}}$ as shown in (27e) and (27d) respectively. Terminal constraints $g_{\mathrm{f},\boldsymbol{\theta}}:\mathcal{S}\rightarrow\mathbb{R}^{n_{g_{\mathrm{f}}}}$ are represented via (27f). To always ensure feasibility, the inequality constraints and terminal constraints need to be relaxed to soft constraints via slack variables $\boldsymbol{\sigma}_{i}\in\mathbb{R}^{n_{g}}$ and $\boldsymbol{\sigma}_{N}\in\mathbb{R}^{n_{g_{\mathrm{f}}}}$. Also, $h_{\boldsymbol{\theta}}$ needs to be designed such there always exists a feasible $\boldsymbol{u}_{i}$. All functions with the index $\boldsymbol{\theta}$ are parameterizable and can be adapted by the RL algorithm. Note that to distinguish the states and actions within the optimal control problem (27) from those within the MDP, we introduce the states and actions within the MPC as $\boldsymbol{x}$ and $\boldsymbol{u}$. After (27) is solved, the MPC policy is derived by taking the first element of the sequence of optimal control actions $\boldsymbol{\mathrm{u}}^{\star}$, so $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})=\boldsymbol{u}_{0}^{\star}$.
@@ 2.3.1 Parametric NLP sensitivities
In order to update the policy parameters with a policy optimization method, the Jacobian of the policy with respect to its parameters needs to be computed. When using MPC as a policy, the derivative of the policy equals the derivative of the solution of the optimization problem, which is only computable explicitly in rare cases. The MPC optimization problem (27) can be categorized as a parametric constrained nonlinear program (NLP). The derivative of the optimal solution with respect to the parameters of the NLP is also often referred to as the (parametric) NLP sensitivities. We show next how the NLP sensitivities can be computed.
@@
$\boldsymbol{z}^{\star}(\boldsymbol{\zeta})=\arg\min_{\boldsymbol{z}}\penalty 10000\ \penalty 10000\$$\Phi(\boldsymbol{z};\boldsymbol{\zeta})$ (28a)
$\mathrm{s.t.}\penalty 10000\ \penalty 10000\$$g(\boldsymbol{z};\boldsymbol{\zeta})\leq 0$ (28b)
@@
where $\boldsymbol{z}\in\mathbb{R}^{n_{\boldsymbol{z}}}$ denotes the decision variables, $\boldsymbol{\zeta}\in\mathbb{R}^{n_{\boldsymbol{\zeta}}}$ denotes the changeable parameters, $\Phi:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}$ denotes the scalar objective function, $g:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}^{n_{g}}$ denotes the inequality constraints, and $h:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}^{n_{h}}$ denotes the equality constraints. The optimal solution $\boldsymbol{z}^{\star}(\boldsymbol{\zeta})$ of (28) changes with the optimization parameters $\boldsymbol{\zeta}$. This functional relationship is typically unknown. We introduce the Lagrangian $\mathcal{L}:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{g}}\times\mathbb{R}^{n_{h}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}$ of (28) as
@@
where $\boldsymbol{\lambda}\in\mathbb{R}^{n_{g}}$ and $\boldsymbol{\nu}\in\mathbb{R}^{n_{h}}$ are the Lagrange multipliers, also known as dual variables, associated to the inequality and equality constraints respectively. To simplify notation, we introduce the stacked primal-dual vector $\boldsymbol{\chi}^{\top}=(\boldsymbol{z}^{\top},\boldsymbol{\lambda}^{\top},\boldsymbol{\nu}^{\top})$ with $\boldsymbol{\chi}\in\mathbb{R}^{n_{\boldsymbol{\chi}}}$ and $n_{\boldsymbol{\chi}}=n_{\boldsymbol{z}}+n_{g}+n_{h}$. Let $\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})$ denote an optimal primal-dual vector solving (28) and satisfying the Karush-Kuhn-Tucker (KKT) conditions (nocedalNumericalOptimization2006), and let $g$ satisfy the linear-independency constraint qualifications (nocedalNumericalOptimization2006), then the equalities of the KKT conditions form an implicit function $F:\mathbb{R}^{n_{\boldsymbol{\chi}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}^{n_{\boldsymbol{\chi}}}$ of the form
@@
where the operator $\odot$ denotes the Hadamard (element-wise) product. In order to obtain the first-order NLP sensitivities of the optimal primal-dual solution with respect to the NLP parameters $\nabla_{\boldsymbol{\zeta}}\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})\in\mathbb{R}^{n_{\boldsymbol{\chi}}\times n_{\boldsymbol{\zeta}}}$, (30) can be implicitly differentiated almost everywhere (fiaccoSensitivityStabilityAnalysis1990) forming the following linear system
@@
with the coefficient matrix $\nabla_{\boldsymbol{\chi}^{\star}}F(\boldsymbol{\chi}^{\star};\boldsymbol{\zeta})\in\mathbb{R}^{n_{\boldsymbol{\chi}}\times n_{\boldsymbol{\chi}}}$ and the right-hand-side matrix $\nabla_{\boldsymbol{\zeta}}F(\boldsymbol{\chi}^{\star};\boldsymbol{\zeta})\in\mathbb{R}^{n_{\boldsymbol{\chi}}\times n_{\boldsymbol{\zeta}}}$. The first-order NLP sensitivities can therefore be obtained by solving (31) for $\nabla_{\boldsymbol{\zeta}}\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})$. The second-order NLP sensitivities can be obtained in a similar fashion and also boil down to the solution of a linear system (brandnerReinforcedModelPredictive2024b; grosReinforcementLearningBased2021). However the right-hand-side matrix requires third partial derivatives of the Lagrangian with respect to full primal-dual vector. This can render the computation intractable, for instance if the number of primal-dual variables $\boldsymbol{\chi}$ is large, or if highly nonlinear terms appear in the NLP such that the third partial derivatives become extremely convoluted.
To compute the deterministic policy gradient (12) and the respective terms of the deterministic policy Hessian (22), NLP (28) must be cast into (27) by setting $\boldsymbol{z}^{\star}=\boldsymbol{\mathrm{u}}^{\star}$ and $\boldsymbol{\zeta}=\boldsymbol{\theta}$. The relevant sensitivities $\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})=\nabla_{\boldsymbol{\theta}}\boldsymbol{u}_{0,\boldsymbol{\theta}}^{\star}(\boldsymbol{s})$ must then be extracted from the full sensitivity matrix.